# Fusing Pretrained ViTs with TCNet for Enhanced EEG Regression

### v1

@@ 1 Introduction
The analysis of Electroencephalogram (EEG) signals is a cornerstone in the advancement of Brain-Computer Interfaces (BCIs), offering profound insights into the intricate neural processes of the human brain. EEG regression, in particular, stands as a pivotal tool in both neuroscience and medical diagnostics, gaining prominence for its ability to decode complex neural dynamics. This technique plays a crucial role in a myriad of applications, ranging from pinpointing brain damage locations to monitoring cognitive activities and deciphering the neural basis of seizures (Subasi and Ercelebi, 2005; Sabbagh et al., 2020; Teplan, 2002). The essence of EEG regression lies in its capacity to transform raw EEG data into interpretable and meaningful information, thus providing an invaluable perspective into the brain’s operations.
@@
1. Innovative Combination of ViTs and TCNet for Advanced EEG Regression: This research marks a significant advancement in EEG regression analysis through the novel integration of pretrained Vision Transformers (ViTs) with Temporal Convolutional Networks (TCNet). This fusion harnesses ViTs’ exceptional capability in processing sequential data and TCNet’s robust feature extraction techniques, culminating in a notable improvement in EEG regression accuracy.In addition, we open source all of our methods and results at github.com/ModeEric/EEGViT-TCNet
@@ 3.3 Training and Evaluation Procedure
Figure 3: An outline of our addition to EEGViT, demonstrating our distinct feature extraction methodology Yang and Modesitt (2023)Bai et al. (2018).
@@ 4.1 Performance Benchmarking
4.2 Computational Efficiency and Speed Enhancement
One of the features of our EEGViT-TCNet model is its computational efficiency. In comparison to the original EEGViT architecture, our enhanced model achieves a substantial speed-up, being 4.32 times faster during inference. This improvement is not merely a testament to the model’s optimized architecture but also underscores its practicality for real-time gaze position prediction applications.
The speed enhancement is achieved through key optimizations in the model’s architecture, mainly through an optimization of the patch size accuracy-efficiency tradeoff. These optimizations ensure that our model not only maintains high accuracy but also operates with greater efficiency, making it well-suited for deployment in computationally expensive environments.
@@ 4.3.1 Impact of Convolutional Layers
We first examined the impact of the additional convolutional layers that bridge the gap between the TCNet and the ViT on the model’s performance. In particular, we analyze the results after removing all possible combinations of 1 convolutional layer (spatial, temporal, and pointwise convolution).
By removing the pointwise layer, we observed a slight increase in the Root Mean Square Error (RMSE) from 51.8 ± 0.6mm to 52.5 ± 0.8mm. This suggests that the pointwise convolutional layer plays a modest role in feature extraction and spatial representation, contributing to the model’s overall accuracy.
For both the spatial and temporal layers we observed a significant increase in the Root Mean Square Error (RMSE) from 51.8 ± 0.6mm to 55.1 ± 0.6mm and 55.0 ± 0.5mm respecitvely. This suggests that both sptial and temproal layers plays a major role in feature extraction and representation, contributing to the model’s overall accuracy.
@@ 5 Discussion
Our novel approach has not only showcased a marked improvement in regression accuracy but also set new benchmarks in processing speed and efficiency. The results obtained from this study reflect the significant potential of leveraging the strengths of both ViTs and TCNet, underscoring the profound impact that such hybrid models can have on understanding and interpreting complex neural dynamics.
The success of the EEGViT-TCNet model in reducing the Root Mean Square Error (RMSE) to its current levels emphasizes the model’s capability to provide a more accurate interpretation of EEG data. This breakthrough is particularly relevant in the development of BCIs, where the precision of signal interpretation directly correlates to the effectiveness and user-friendliness of the interface. In clinical settings, the enhanced accuracy and speed of EEG analysis facilitated by the EEGViT-TCNet model could lead to more timely and accurate diagnoses of neurological conditions, potentially transforming patient care.
@@
Moreover, the exploration of other hybrid architectures and their efficacy across various domains of neural data presents an exciting avenue for research. The integration of multimodal data sources, alongside the application of transfer learning techniques, could further refine the accuracy and applicability of EEG signal analysis methods. Such advancements could pave the way for the development of more sophisticated BCIs, offering improved interaction mechanisms between humans and machines.
@@ 6 Conclusion
This study represents a significant advancement in EEG regression analysis, underscoring the role of meticulous feature extraction in the efficacy of sophisticated computational models like Vision Transformers (ViT). The integration of Temporal Convolutional Networks (TCNet) with pretrained ViTs has unveiled the vast potential of harmonizing specialized feature extraction techniques with advanced deep learning frameworks. This synergy not only elevates the accuracy of EEG analysis but also establishes a new benchmark in the field, showcasing the profound benefits of refined feature representation.
@@
For future research, the horizon of EEG analysis and deep learning promises continued expansion and innovation. The development of increasingly sophisticated models capable of navigating the complexities inherent in EEG data is anticipated. A pivotal challenge will be enhancing the interpretability of these models, ensuring they not only perform optimally but also offer actionable insights for practitioners. For example, the performance of nearly all effective models in EEG analysis currently relies heavily on the subject-dependent nature on which they are trained, leading to the possibility of poor generalization in some instances. Moreover, integrating these advanced models into real-world applications will be crucial, extending the benefits of this research to society at large.

### v2

@@ Abstract
Keywords:
EEG Analysis Temporal Convolutional Networks Brain Computer Interfaces Vision Transformers
@@ 1 Introduction
Analyzing Electroencephalogram (EEG) signals is fundamental to the progress of Brain-Computer Interfaces (BCIs), offering deep insights into the complex neural processes of the human brain. In the past decade, a wide range of machine learning and deep learning algorithms have been applied to EEG data, resulting in significant advancements in various applications. These applications encompass emotion recognition, motor imagery, mental workload evaluation, seizure detection, Alzheimer’s disease classification, sleep stage scoring, and many others (Craik et al., 2019; Roy et al., 2019; Altaheri et al., 2023; Qu, 2022; Gao et al., 2021; Hossain et al., 2023; Yi and Qu, 2022; Key et al., 2024; Li et al., 2024; Koome Murungi et al., 2023; Dou et al., 2022; Zhou et al., 2022; Qu et al., 2020b, c, a, 2018, 2019; Murungi et al., 2023; Saeidi et al., 2021; Qu and Hickey, 2022; Rasheed et al., 2020; Wang and Qu, 2022; Dadebayev et al., 2022; Li et al., 2020; Aggarwal and Chugh, 2022). EEG regression, in particular, stands as a pivotal tool in both neuroscience and medical diagnostics, gaining prominence for its ability to decode complex neural dynamics. This technique plays a crucial role in a myriad of applications, ranging from pinpointing brain damage locations to monitoring cognitive activities and deciphering the neural basis of seizures (Subasi and Ercelebi, 2005; Sabbagh et al., 2020; Teplan, 2002). The essence of EEG regression lies in its capacity to transform raw EEG data into interpretable and meaningful information, thus providing an invaluable perspective into the brain’s operations.
@@
1. Innovative Combination of ViTs and TCNet for Advanced EEG Regression: This research marks a significant advancement in EEG regression analysis through the novel integration of pretrained Vision Transformers (ViTs) with Temporal Convolutional Networks (TCNet). This fusion harnesses ViTs’ exceptional capability in processing sequential data and TCNet’s robust feature extraction techniques, culminating in a notable improvement in EEG regression accuracy.
@@ 3.3 Training and Evaluation Procedure
Figure 3: An outline of our addition to EEGViT, demonstrating our distinct feature extraction methodology.
@@ 4.2.1 Impact of Convolutional Layers
We first examined the impact of the additional convolutional layers that bridge the gap between the TCNet and the ViT on the model’s performance. In particular, we analyze the results after removing all possible combinations of 1 convolutional layer (spatial, temporal, and pointwise convolution). By removing the pointwise layer, we observed a slight increase in the Root Mean Square Error (RMSE) from 51.8 ± 0.6mm to 52.5 ± 0.8mm. This suggests that the pointwise convolutional layer plays a modest yet significant role in feature extraction and spatial representation, contributing to the model’s overall accuracy.
@@ 5 Discussion
The integration of Vision Transformers (ViTs) with Temporal Convolutional Networks (TCNet) within the EEGViT-TCNet model represents a substantial leap forward in EEG signal analysis, particularly for applications within brain-computer interfaces (BCIs) and the broader realm of neural signal processing. This novel approach has not only showcased a marked improvement in regression accuracy but also set new precedents in processing speed and efficiency. The results obtained from this study reflect the significant potential of leveraging the strengths of both ViTs and TCNet, underscoring the profound impact that such hybrid models can have on understanding and interpreting complex neural dynamics.
The success of the EEGViT-TCNet model in reducing the Root Mean Square Error (RMSE) to unprecedented levels emphasizes the model’s capability to provide a more accurate interpretation of EEG data. This breakthrough is particularly relevant in the development of BCIs, where the precision of signal interpretation directly correlates to the effectiveness and user-friendliness of the interface. In clinical settings, the enhanced accuracy and speed of EEG analysis facilitated by the EEGViT-TCNet model could lead to more timely and accurate diagnoses of neurological conditions, potentially transforming patient care.
@@
Moreover, the exploration into other hybrid architectures and their efficacy across various domains of neural data presents an exciting avenue for research. The integration of multimodal data sources, alongside the application of transfer learning techniques, could further refine the accuracy and applicability of EEG signal analysis methods. Such advancements could pave the way for the development of more sophisticated BCIs, offering improved interaction mechanisms between humans and machines.
@@ 6 Conclusion
This study represents a significant advancement in EEG regression analysis, underscoring the indispensable role of meticulous feature extraction in the efficacy of sophisticated computational models like Vision Transformers (ViT). The integration of Temporal Convolutional Networks (TCNet) with pretrained ViTs has unveiled the vast potential of harmonizing specialized feature extraction techniques with advanced deep learning frameworks. This synergy not only elevates the accuracy of EEG analysis but also establishes a new benchmark in the field, showcasing the profound benefits of refined feature representation.
@@
As we chart the course for future research, the horizon of EEG analysis and deep learning promises continued expansion and innovation. The development of increasingly sophisticated models capable of navigating the complexities inherent in EEG data is anticipated. A pivotal challenge will be enhancing the interpretability of these models, ensuring they not only perform optimally but also offer actionable insights for practitioners. Moreover, integrating these advanced models into real-world applications will be crucial, extending the benefits of this research to society at large. Additionally, exploring various deep learning techniques on different datasets for comparative studies (An et al. (2023a, b); Jiang et al. (2023); Gui et al. (2024); Lu et al. (2023); Chen et al. (2024); Ma (2022); Ma et al. (2024); Tan et al. (2023, 2021); Qiu et al. (2023); Zhao et al. (2024); Zhang et al. (2022, 2023)) could provide valuable insights and further enhance the field.